OK, I am not a student, but I have recently been asked to find an answer for this equation by a friend. Everything I can do, or find on line, says that if the values are different numbers, there is no viable solution. Please help me with this.
X*Y = X+Y
What are the values of X and Y? Or is there no actual answer? Thank you.

Answer by stanbon(75887)

X*Y = X+Y
What are the values of X and Y?
---------------
xy=x+y
x+y-xy=0
y-xy=-x
y(1-x)=-x
y=-x/(1-x)
----------
x may take on any value except x=1
----
For example:
If x=3, then y=-3/-2=3/2
So the pair x=3, y=3/2 are two different numbers that
satisfy the equation.
There are an infinite number of solutions to your question.
Cheers,
Stan H.